# Machine Space I: Weak exponentials and quantification over compact spaces

### v4

@@ 4. Compactness and universal quantification
We believe that machine space also has the potential to be useful in clarifying other aspects of topology. In this section we use it to study compactness. Escardó has explained compactness in terms of an algorithm for universal quantification in the setting of semantics of programming languages [7, 9, 10]. Our approach yields a version of this algorithm that involves only topological/localic concepts and the core operations provided by the verifiability interpretation, without restricting us to a particular programming language. Thus, we give a positive answer to second question posed in the introduction: there is a uniform algorithm for universal quantification over any compact space.
@@ 4.1. An algorithm for universal quantification
It is important to understand exactly what it is such an algorithm needs to do. Of course, to give such a procedure we need to know a precise description of the space $X$ by generators and relations. From this we can mathematically derive whether a particular formal combination of generators is equal to $1$. On the other hand, recall that ‘the exact composition’ of the machines in machine space is opaque to us. Indeed, the opens of machine space only allow us to test very particular properties of the machines. In the first case we are dealing with discrete ‘syntax’, while in the second we are working with abstract machines which we do not have knowledge about a priori (see Remark˜3.6).
@@
Note that the algorithm itself involves testing whether $\bigvee_{F\in S}\bigwedge_{g\in F}g$ is a cover of $X$. This is on the ‘axiomatic’ level of Remark˜3.6 and how to test this is determined mathematically before running the algorithm. The role of the algorithm is to translate from the axiomatic level to the level of opaque machines.
Theorem 4.1.
Let $X$ be a compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. Algorithm˜1 semi-decides if the given machine $m$ halts on all of $X$.
Proof.
@@
Conversely, if the computation halts on $m$. Then there is some $S$ which provides a finite refinement of the cover given by $m$ which covers $X$. Hence $\bigvee_{i\in I}\bigwedge_{j\in J_{i}}g_{j}$ covers $X$ and $m$ halts on all of $X$, as required. ∎
Remark 4.2.
@@
In ˜3.7 the procedure associated to ${\boxtimes}F$ is simply the act of observing the machine and comparing the generators it visits to the finite collection $F$. It is clear that this can be done uniformly in $F$. Alternatively, we might regard a machine as a program $m\colon\Sigma^{G}\to\Sigma$. To test whether a branch is contained in a subset $F$, we simply supply as input the generalised point $p_{F}\colon G\to\Sigma$ which halts precisely on the elements of $F$.
Example 4.3 (Cantor space).
Cantor space is the space $2^{\mathbb{N}}$ of infinite binary sequences. A presentation is given by $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1\rangle$. Intuitively, $z_{n}$ should be thought of as the open consisting of the sequences whose $n^{\text{th}}$ digit is 0 and $u_{n}$ the open of sequences with $1$ in the $n^{\text{th}}$ position. A machine will be given by a formal expression of the form $\bigvee_{i\in I}\left(\bigwedge_{j\in J_{i}}z_{j}\wedge\bigwedge_{k\in K_{i}}u_{k}\right)$.
@@
The implementation of $\mathtt{test}(m\in{\boxtimes}F)$ depends on the precise model. In any case, we arrive at an algorithm which achieves the same ends as that given by Escardó in [7, 10]. ∎
We will explore how our approach relates to other representations of spaces and the link to Escardó’s setting in Section˜4.3.
Example 4.4 (The closed interval).
Consider the case of the closed real interval $X=[0,1]$. We can take a generating set consisting of certain intervals with rational end points, $G=\{[0,q)\mid q\in\mathbb{Q}\cap(0,1]\}\cup\{(q,1]\mid q\in\mathbb{Q}\cap[0,1)\}$. Then in the computational model, $[0,1]$ embeds as a subset of partial functions on $G$ and so a real $r$ corresponds to a function that halts on $[0,q)$ or $(q,1]$ if and only if $r\in[0,q)$ or $r\in(q,1]$ respectively. Then for a ‘machine’ $m\colon\Sigma^{G}\to\Sigma$ the algorithm halts if $m$ halts on all of the (functions corresponding to) reals in $[0,1]$. It is elementary to compute when a formal join of rational intervals covers $[0,1]$. Finally, $\mathtt{test}(m\in{\boxtimes}F\})$ is given by simply calling $m(p_{F})$ where $p_{F}\in\Sigma^{G}$ is given by $p_{F}(g)=\top\iff g\in F$. ∎
@@
Remark 4.5.
Constructively, there is a ‘dual’ notion to compactness called overtness (see [26, 15]) which has the same relation to existential quantification as compactness has to universal quantification. One can give an algorithm for existential quantification that is very similar to Algorithm˜1 except that instead of iterating over covers we iterate over ‘positive’ (classically: non-zero) elements. Classically, every locale is overt, though the resulting algorithm is still more subtle than naively searching through every real number in turn. Instead, it would use the fact that non-trivial open intervals with rational endpoints form a countable base for $\mathbb{R}$, which has the advantage of avoiding uncountable parallelism.
@@ 4.2. Topological consequences
It is possible to unwind Algorithm˜1 to give an open of $\Sigma^{\Sigma^{G}}$. Each expression $\mathtt{test}(m\in{\boxtimes}F)$ is replaced with the corresponding open ${\boxtimes}F$. The second for loop takes the meet of these opens, $\bigwedge_{F\in S}{\boxtimes}F$. The first for (together with the if) corresponds to taking the join over all $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ such that $\bigvee_{F\in S}\bigwedge_{g\in F}g$ covers $X$. Let us denote the set of such $S$ by $\mathcal{C}$. Thus, the resulting open is $\bigvee_{S\in\mathcal{C}}\bigwedge_{F\in S}{\boxtimes}F$. When $X$ is compact, this is the open of machines which cover $X$. The following result is then immediate from Theorem˜4.1.
Corollary 4.6.
@@
Also note Corollary˜4.6 easily generalises to arbitrary compact sublocales $K$ of a general locale $X$ by simply viewing $K$ as direct sublocale of $\Sigma^{G}$.
Remark 4.7.
Corollary˜4.6 also follows from the Hofmann–Mislove theorem [12]. Moreover, the Hofmann–Mislove theorem also provides a converse giving that the set of all covering machine of $X$ is Scott-open if and only if $X$ is compact. (See also [7, Lemma 7.4].) However, we believe restriction to machine space and the explicit algorithm provide an interesting new perspective on the situation.
Remark 4.8.
Of course, in the overt case we can similarly extract an open of all ‘positive’ machines for $X$ from the algorithm mentioned in Remark˜4.5.
@@ 4.3. Link to domain-theoretic approaches
Escardó considers the data type of infinite binary sequences. We imagine these are implemented as streams where the terms are computed in order one by one on demand. Such data consists not only of infinite sequences, but also additional elements corresponding to sequences whose first $n$ terms are defined, but for which accessing later terms in the sequence causes the computation to hang.
@@
Example 4.9.
Let us consider the case of Cantor space from Example˜4.3. Recall that $\mathcal{O}(2^{\mathbb{N}})$ can be expressed as a quotient of $\mathcal{O}(\Sigma^{G})$ where $G=\{z_{n}\mid n\in\mathbb{N}\}\sqcup\{u_{n}\mid n\in\mathbb{N}\}$ by the congruence $C=\langle z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1,\,n\in\mathbb{N}\rangle$. The closure of such a sublocale corresponds to the congruence generated by the pairs $(0,c)\in C$. In this case we obtain the following presentation for the closure: $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0\rangle$. This describes a dcpo the elements of which can be thought of as partial functions from $\mathbb{N}$ to $2=\{0,1\}$. The points of Cantor space correspond to the total functions, which do appear as the maximal elements of this dcpo. ∎
In general, the compactness algorithm works equally well if we replace $\mathcal{P}(G)$ with such a continuous dcpo $D$, or indeed any locally compact sublocale of machine space which contains $X$. The algorithm takes an element $m$ of $\Sigma^{D}$ as input and simply applies Algorithm˜1 to $s(m)$ where $s$ is a section of the quotient $\Sigma^{\Sigma^{G}}\to\Sigma^{D}$ from Proposition˜3.10.
@@
The above representation of Cantor space is not quite the one used by Escardó, since he views Cantor space not as a function, but as a stream — that is, there is an ordering on indices so that if the value at $i$ is defined, so are all previous values. This is related to alternative presentation of Cantor space in terms of prefixes.
Example 4.10.
A presentation for $2^{\mathbb{N}}$ can be given with generators $\ell_{p}$ corresponding to sequences with finite prefix $p\in 2^{*}$. We have
@@
However, taking the closure of this sublocale of $\Sigma^{2^{*}}$ does not yield the correct dcpo. This is because it ignores the natural order structure on the prefixes. The solution is to modify the parent space $\Sigma^{G}$, so that instead of taking the free frame on a set $G$, we take the free frame on a poset. In this case, we equip $2^{*}$ with the reverse of the prefix order. Then $\mathcal{O}2^{\mathbb{N}}$ is the quotient of this frame by the congruence generated by the second and third relations above (since the first relation is already handled by the order structure on $G$). The closure of this sublocale then has the presentation $\langle\{\ell_{p}:p\in(2^{*})^{\mathrm{op}}\}\in\mathrm{Pos}\mid\ell_{p}\wedge\ell_{q}=0\text{ if $q\nprec p$ and $p\nprec q$}\rangle$. It is not hard to see that this is the locale of finite and infinite sequences binary sequences. (Indeed, it is the third relation in the presentation for $2^{\mathbb{N}}$ above that forces all sequences to be infinite.) This is precisely the dcpo considered by Escardó. (Finite sequences are interpreted as sequences which hang after some point.)

### v5

@@ 4. Compactness and universal quantification
We believe that machine space also has the potential to be useful in clarifying other aspects of topology. In this section we use it to study compactness. Escardó has explained compactness in terms of an algorithm for universal quantification in the setting of semantics of programming languages [7, 9, 10]. Our approach yields a version of this algorithm that involves only topological/localic concepts and the core operations provided by the verifiability interpretation, without restricting us to a particular programming language. Thus, we give a positive answer to the second question posed in the introduction: there is a uniform algorithm for universal quantification over any compact space.
@@ 4.1. An algorithm for universal quantification
It is important to understand exactly what it is such an algorithm needs to do. Of course, to give such a procedure we need to know a precise description of the space $X$ by generators and relations. From this we can mathematically derive whether a particular formal combination of generators is equal to $1$. On the other hand, recall that ‘the exact composition’ of the machines in machine space is opaque to us. Indeed, the opens of machine space only allow us to test very particular properties of the machines. In the first case we are dealing with discrete ‘syntax’, while in the second we are working with abstract machines which we do not have knowledge about a priori (see 3).
@@
Note that the algorithm itself involves testing whether $\bigvee_{F\in S}\bigwedge_{g\in F}g$ is a cover of $X$. This is on the ‘axiomatic’ level of 3 and how to test this is determined mathematically before running the algorithm. The role of the algorithm is to translate from the axiomatic level to the level of opaque machines.
Theorem 13.
Let $X$ be a compact locale with $\mathcal{O}X=\langle G\mid R\rangle$. Algorithm 1 semi-decides if the given machine $m$ halts on all of $X$.
Proof 4.1.
@@
Conversely, if the computation halts on $m$, then there is some $S$ which provides a finite refinement of the cover given by $m$ which covers $X$. Hence $\bigvee_{i\in I}\bigwedge_{j\in J_{i}}g_{j}$ covers $X$ and $m$ halts on all of $X$, as required.
Remark 14.
@@
In Model 4 the procedure associated to ${\boxtimes}F$ is simply the act of observing the machine and comparing the generators it visits to the finite collection $F$. It is clear that this can be done uniformly in $F$. Alternatively, we might regard a machine as a program $m\colon\Sigma^{G}\to\Sigma$. To test whether a branch is contained in a subset $F$, we simply supply as input the generalised point $p_{F}\colon G\to\Sigma$ which halts precisely on the elements of $F$.
{exa}
[Cantor space] Cantor space is the space $2^{\mathbb{N}}$ of infinite binary sequences. A presentation is given by $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1\rangle$. Intuitively, $z_{n}$ should be thought of as the open consisting of the sequences whose $n^{\text{th}}$ digit is 0 and $u_{n}$ the open of sequences with $1$ in the $n^{\text{th}}$ position. A machine will be given by a formal expression of the form $\bigvee_{i\in I}\left(\bigwedge_{j\in J_{i}}z_{j}\wedge\bigwedge_{k\in K_{i}}u_{k}\right)$.
@@
The implementation of $\mathtt{test}(m\in{\boxtimes}F)$ depends on the precise model. In any case, we arrive at an algorithm which achieves the same ends as that given by Escardó in [7, 10]. ∎ We will explore how our approach relates to other representations of spaces and the link to Escardó’s setting in subsection 4.3.
{exa}
[The closed interval] Consider the case of the closed real interval $X=[0,1]$. We can take a generating set consisting of certain intervals with rational end points, $G=\{[0,q)\mid q\in\mathbb{Q}\cap(0,1]\}\cup\{(q,1]\mid q\in\mathbb{Q}\cap[0,1)\}$. Then in the computational model, $[0,1]$ embeds as a subset of partial functions on $G$ and so a real $r$ corresponds to a function that halts on $[0,q)$ or $(q,1]$ if and only if $r\in[0,q)$ or $r\in(q,1]$ respectively. Then for a ‘machine’ $m\colon\Sigma^{G}\to\Sigma$ the algorithm halts if $m$ halts on all of the (functions corresponding to) reals in $[0,1]$. It is elementary to compute when a formal join of rational intervals covers $[0,1]$. Finally, $\mathtt{test}(m\in{\boxtimes}F\})$ is given by simply calling $m(p_{F})$ where $p_{F}\in\Sigma^{G}$ is given by $p_{F}(g)=\top\iff g\in F$. ∎
@@
Remark 15.
Constructively, there is a ‘dual’ notion to compactness called overtness (see [26, 15]) which has the same relation to existential quantification as compactness has to universal quantification. One can give an algorithm for existential quantification that is very similar to Algorithm 1 except that instead of iterating over covers we iterate over ‘positive’ (classically: non-zero) elements. Classically, every locale is overt, though the resulting algorithm is still more subtle than naively searching through every real number in turn. Instead, it would use the fact that non-trivial open intervals with rational endpoints form a countable base for $\mathbb{R}$, which has the advantage of avoiding uncountable parallelism.
@@ 4.2. Topological consequences
It is possible to unwind Algorithm 1 to give an open of $\Sigma^{\Sigma^{G}}$. Each expression $\mathtt{test}(m\in{\boxtimes}F)$ is replaced with the corresponding open ${\boxtimes}F$. The second for loop takes the meet of these opens, $\bigwedge_{F\in S}{\boxtimes}F$. The first for (together with the if) corresponds to taking the join over all $S\in\mathcal{P}_{\mathsf{fin}}(\mathcal{P}_{\mathsf{fin}}(G))$ such that $\bigvee_{F\in S}\bigwedge_{g\in F}g$ covers $X$. Let us denote the set of such $S$ by $\mathcal{C}$. Thus, the resulting open is $\bigvee_{S\in\mathcal{C}}\bigwedge_{F\in S}{\boxtimes}F$. When $X$ is compact, this is the open of machines which cover $X$. The following result is then immediate from 13.
Corollary 16.
@@
Also note 16 easily generalises to arbitrary compact sublocales $K$ of a general locale $X$ by simply viewing $K$ as direct sublocale of $\Sigma^{G}$.
Remark 17.
16 also follows from the Hofmann–Mislove theorem [12]. Moreover, the Hofmann–Mislove theorem also provides a converse giving that the set of all covering machines of $X$ is Scott-open if and only if $X$ is compact. (See also [7, Lemma 7.4].) However, we believe restriction to machine space and the explicit algorithm provide an interesting new perspective on the situation.
Remark 18.
Of course, in the overt case we can similarly extract an open of all ‘positive’ machines for $X$ from the algorithm mentioned in 15.
@@ 4.3. Link to domain-theoretic approaches
Escardó considers the data type of infinite binary sequences. We imagine these are implemented as streams where the terms are computed in order one by one on demand. Such data consists not only of infinite sequences, but also additional elements corresponding to sequences whose first $n$ terms are defined, but for which accessing later terms in the sequence causes the computation to diverge.
@@
{exa}
Let us consider the case of Cantor space from 14. Recall that $\mathcal{O}(2^{\mathbb{N}})$ can be expressed as a quotient of $\mathcal{O}(\Sigma^{G})$ where $G=\{z_{n}\mid n\in\mathbb{N}\}\sqcup\{u_{n}\mid n\in\mathbb{N}\}$ by the congruence $C=\langle z_{n}\wedge u_{n}=0,z_{n}\vee u_{n}=1,\,n\in\mathbb{N}\rangle$. The closure of such a sublocale corresponds to the congruence generated by the pairs $(0,c)\in C$. In this case we obtain the following presentation for the closure: $\langle z_{n},u_{n},\,n\in\mathbb{N}\mid z_{n}\wedge u_{n}=0\rangle$. This describes a dcpo the elements of which can be thought of as partial functions from $\mathbb{N}$ to $2=\{0,1\}$. The points of Cantor space correspond to the total functions, which do appear as the maximal elements of this dcpo. ∎
In general, the compactness algorithm works equally well if we replace $\mathcal{P}(G)$ with such a continuous dcpo $D$, or indeed any locally compact sublocale of machine space which contains $X$. The algorithm takes an element $m$ of $\Sigma^{D}$ as input and simply applies Algorithm 1 to $s(m)$ where $s$ is a section of the quotient $\Sigma^{\Sigma^{G}}\to\Sigma^{D}$ from 6.
@@
The above representation of Cantor space is not quite the one used by Escardó, since he views Cantor space not as a function, but as a stream — that is, there is an ordering on indices so that if the value at $i$ is defined, so are all previous values. This is related to an alternative presentation of Cantor space in terms of prefixes. {exa} A presentation for $2^{\mathbb{N}}$ can be given with generators $\ell_{p}$ corresponding to sequences with finite non-empty prefix $p\in 2^{+}\coloneqq 2^{*}\setminus\{\varepsilon\}$. We have
@@
However, taking the closure of this sublocale of $\Sigma^{2^{+}}$ does not yield the correct dcpo. This is because it ignores the natural order structure on the prefixes. The solution is to modify the parent space $\Sigma^{G}$, so that instead of taking the free frame on a set $G$, we take the free frame on a poset. In this case, we equip $2^{+}$ with the reverse of the prefix order. Then $\mathcal{O}2^{\mathbb{N}}$ is the quotient of this frame by the congruence generated by the second, third and fourth relations above (since the first relation is already handled by the order structure on $G$). The closure of this sublocale has the presentation $\langle\{\ell_{p}:p\in(2^{+})^{\mathrm{op}}\}\in\mathrm{Pos}\mid\ell_{p}\wedge\ell_{q}=0\text{ if $q\nprec p$ and $p\nprec q$}\rangle$. It is not hard to see that this is the locale of finite and infinite binary sequences. (Indeed, it is the final two relations in the presentation for $2^{\mathbb{N}}$ above that forces all sequences to be infinite.) This is precisely the dcpo considered by Escardó. (Finite sequences are interpreted as sequences which are undefined after some point.)
@@
Acknowledgements
We thank the referees for their comments, which improved the presentation of this paper.